Center Sports is considering a new automated maintenance machine that will lower costs by $330,161 per year. The machine costs $797,987 and will be depreciated straight line to zero over the 8 year life of the equipment. If Center is taxed at a rate of 12%, what is the operating cash flow associated with the purchase? (Round your answer to the nearest dollar)
Answer #1

Annual depreciation=(Cost-Salvage Value)/Useful Life

=797,987/8=$99748.375

OCF=Reduction in costs(1-tax rate)+Tax savings on Annual depreciation

=330,161*(1-0.12)+(99748.375*0.12)

=$302511(Approx).